JamesTDG:

I need somebody to help explain this question to me: (SC will be posted below) The volume of a right triangular prism is 72 cubic feet. The height of the prism is 9 feet. The triangular base is an isosceles right triangle. A right triangular prism has a volume of 72 cubic feet. Triangles A E C and B F D are the base triangles. Sides F D and B F are congruent and form a right angle. The height of the prism is 9 feet. What is the area of the triangular base? 16 square feet What is the length of edge DF? 8 feet

JamesTDG:

(I left guesses)

JamesTDG:

What I have to go off of: The height is 9, which would be the tallest point on the triangular prism, which is BF. The image shows that BF=DF, so DF should also be 9 meaning that the area should be (9*9)/2, but it is not that

axie:

@snowflake0531

snowflake0531:

Volume is base times height, so base times 9 equals 72 So what's base?

JamesTDG:

CDEF is the base

snowflake0531:

@jamestdg wrote:
CDEF is the base
no it's not

JamesTDG:

DBF?

snowflake0531:

An easy way to find is to imagine... what 2-D shape stacked up together will create that prism, in this case, it's the triangle, so yes DBF

JamesTDG:

What I was given says that the height is 9, and since bf is the tallest point it is 9

JamesTDG:

right?

snowflake0531:

...no

snowflake0531:

Since you already know that base is DBF... height would be AB/CD/EF, (they're all the same length)

JamesTDG:

ok

snowflake0531:

So what's the area of the base if base times 9 is 72

JamesTDG:

plugging 9*x=72 in my graphing calculator says 8

snowflake0531:

damn, you used a calculator but yes, it's 8 So now \[\frac{x^2}{2}=8\\x^2=16\] so what's the x

JamesTDG:

solve the square root of 16 to get 4, x would then be 4

snowflake0531:

yes

snowflake0531:

so do you get your answers now?

JamesTDG:

yep, thanks

snowflake0531:

yw~